OpenStudy (anonymous):

Find the LCD that would eliminate the fractions in this equation. X/2+2/3=1/4

OpenStudy (anonymous):

\[\frac{ x }{ 2 }+\frac{ 2 }{ 3 }-\frac{ 1 }{ 4 }=0\] \[\frac{ x }{ 2 }+\frac{ 8 }{ 12 }-\frac{ 3 }{ 12 }=0\] \[\frac{ 1x }{ 2 }+\frac{ 5 }{ 12 }=0\] LCD IS 12

OpenStudy (anonymous):

i am ganna post another one in a little bit